Calculate the number of pounds of CO2 released into the atmosphere when a 16.0 gallon tank of gasoline is burned in an automobile engine. Assume that gasoline is primarily octane, C8H18, and that the density of gasoline is 0.692 g⋅mL−1. This assumption ignores additives. Also, assume complete combustion.
Useful conversion factors:
1 gallon=3.785 L
1 kg=2.204 lb
1
Expert's answer
2019-10-04T06:46:15-0400

Combustion of octane is given by:

"2C_8H_{18}+25O_2\to 16CO_2+18H_2O"

Molar mass of octane = "114.23\ g\ mol^{-1}"

228.46 gm of octane will produce 704 gm of carbondioxide


Mass of octane taken = "692\times 16\times 3.785=41907.52\ g"

41907.52 gm octane will produce "\frac{704\times41907.52}{228.46}=129,138.12\ gm\ CO_2"

= "129.13812\ kg\ CO_2"

"129.13812\times 2.204\ lb"

"=284.6\ pounds\ CO_2"